# Decision-focused predictions via pessimistic bilevel optimization: complexity and algorithms

### v4

@@ 2 Complexity results
$\hat{u}_{k}=\left\{\begin{array}[]{ll}1&\text{if }(\omega_{u})_{k}\leq 0\\ 0&\text{if }(\omega_{u})_{k}>0\end{array}\right.$
and accordingly, set
$\hat{y}_{j}=\left\{\begin{array}[]{ll}1&\text{ if }f_{j}(u_{j_{\ell}})=1\text{for some $\ell=1,2,3$}\\ 0&\text{ otherwise}\end{array}\right.$
Note that $\hat{y}_{j}$ is 1 if and only if the assignment $\hat{u}$ satisfies clause $C_{j}$. It is not hard to see that $(\hat{u},\hat{y})$ satisfies (7b) (i.e., it is optimal for the lower-level problem), and thus it is feasible for the maximization problem. But since we know the maximum is at most $-m$, we must have
$-\sum_{j=1}^{m}\hat{y}_{j}\leq-m.$
This means there are at least $m$ clauses satisfied, i.e., the 3SAT formula is satisfiable.
Theorem 2.2
[…]
Proof
[…]
$\max\quad$ $\frac{1}{N}\sum_{i=1}^{N}(c^{i})^{\top}v^{i}$
s.t $v^{i}\in\arg\min\,\big{\{}(x^{i}\hat{\omega})^{\top}v\,:\,Av\leq b\big{\}}$
$\max\quad$ $(\rho^{i})^{\top}b$
s.t $(\rho^{i})^{\top}A=x^{i}\hat{\omega}$
$\rho^{i}\leq 0.$
For each $i$, let us consider $\hat{\rho}^{i}$ in the relative interior of the optimal face of the dual. This means that the optimal face of the primal can be described as
$F^{i}(\hat{\omega}):=\{v\in V\,:\,a_{j}^{\top}v=b_{j},\,j\,:\,\hat{\rho}^{i}_{j}<0\}$
We claim that any $(\tilde{\omega},\tilde{\rho})$ satisfying
$(\tilde{\rho}^{i})^{\top}A$ $=x^{i}\tilde{\omega}$ $\forall i\in[N]$ (8a)
$\tilde{\rho}^{i}_{j}$ $\leq-1$ $\hat{\rho}^{i}_{j}<0,\,\forall i\in[N]$ (8b)
$\tilde{\rho}^{i}_{j}$ $=0$ $\hat{\rho}^{i}_{j}=0,\,\forall i\in[N]$ (8c)
is such that $\arg\min\{(x^{i}\tilde{\omega})^{\top}v\,:\,Av\leq b\}=F^{i}(\hat{\omega})$ for every $i\in[N]$. Indeed, if there are such $\tilde{\omega}$ and $\tilde{\rho}$, then
$\arg\min\{(x^{i}\tilde{\omega})^{\top}v\,:\,Av\leq b\}$
$=$ $\arg\min\{(\tilde{\rho}^{i})^{\top}Av\,:\,Av\leq b\}$
$=$ $\arg\min\left\{\sum_{j:\hat{\rho}^{i}_{j}<0}\tilde{\rho}^{i}_{j}a_{j}^{\top}v\,:\,Av\leq b\right\}$
The last objective function is lower bounded by $\sum_{j:\hat{\rho}^{i}_{j}<0}\tilde{\rho}^{i}_{j}b_{j}$ (recall that the $\tilde{\rho}$ values are non-positive). Moreover, this lower bound is met if and only if
$a_{j}^{\top}v=b_{j}\qquad\forall j\,:\,\hat{\rho}^{i}_{j}<0.$
Therefore, we obtain that each $\tilde{\rho}^{i}$ is dual optimal for $\min\{(x^{i}\tilde{\omega})^{\top}v\,:\,Av\leq b\}$ and $\arg\min\{(x^{i}\tilde{\omega})^{\top}v\,:\,Av\leq b\}=F^{i}(\hat{\omega})$.
To conclude, note that (8) is always feasible, as a rescaled version of $(\hat{\omega},\hat{\rho})$ satisfies the system. Since the coefficients of (8) are given by the entries of $A,x^{i}$, zeros, and ones, we conclude that there must be a $(\tilde{\omega},\tilde{\rho})$ of polynomial size that satisfies the system.
Corollary 1
[…]
@@ 3 A non-convex quadratic reformulation
While the previous section establishes that finding a worst-case efficient method for solving (4) is unlikely, we can still hope to find methods with good practical performance.
In this section, we will apply duality arguments, supported by the assumption that $V$ in problem (1) is non-empty and bounded, in order to obtain a more manageable formulation of the problem.
Under our assumptions, we can assume (1) has the following form:
$\min_{v}\quad$ $c^{\top}v$ (9a)
s.t. $Av\geq b$ (9b)
$v\leq 1$ (9c)
$v\geq 0.$ (9d)
@@
[…]
$\quad\quad\quad\quad\quad\quad\quad\tilde{v}^{i}\leq 1$
$\quad\quad\quad\quad\quad\quad\quad\tilde{v}^{i}\geq 0.$
@@ 3.1 Duality arguments
A common approach to solving optimistic bilevel problems involving convex lower-level problems is to reformulate them by replacing the lower-level problems with their optimality (Karush-Kuhn-Tucker or KKT) conditions (see Kleinert et al., (2021)). In this subsection, we follow the same type of argument twice to achieve a single-level reformulation of the pessimistic bilevel problem (10).
@@
$\quad v^{i}\leq 1$ $\forall i\in[N]$ (11c)
$\quad v^{i}\geq 0$ $\forall i\in[N]$ (11d)
$\quad A^{\top}\rho^{i}+\alpha^{i}\leq m(\omega,x^{i})$ $\forall i\in[N]$ (11e)
$\quad\rho^{i}\geq 0$ $\forall i\in[N]$ (11f)
$\quad\alpha^{i}\leq 0$ $\forall i\in[N]$ (11g)
$\quad m(\omega,x^{i})^{\top}v^{i}\leq b^{\top}\rho^{i}+\mathbf{1}^{\top}\alpha^{i}$ $\forall i\in[N]$ (11h)
In this formulation (11b)-(11d) impose primal feasibility, (11e)-(11g) impose dual feasibility, and (11h) imposes strong duality [^1] Strong duality is typically written as an equality constraint; however, the $\geq$ inequality always holds due to weak duality..
@@
$\min_{\omega,\mu,\theta,\delta,\gamma}\,$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+\mathbf{1}^{\top}\theta^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)$ (12a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\gamma^{i}+\theta^{i}\geq\frac{1}{N}c^{i}$ $\forall i\in[N]$ (12b)
@@
$-\gamma^{i}\mathbf{1}+\delta^{i}\leq 0$ $\forall i\in[N]$ (12d)
$\mu^{i}\leq 0$ $\forall i\in[N]$ (12e)
$\delta^{i},\gamma^{i},\theta^{i}\geq 0$ $\forall i\in[N]$ (12f)
@@ 3.2 Shortest path as a bounded linear program
In this work, and motivated by Elmachtoub and Grigas, (2022), we consider the shortest path problem with unknown cost vectors. It is well known that this problem can be formulated as a linear program using a totally unimodular constraint matrix. However, the feasible region may not be bounded, as the underlying graph may have negative cycles, which correspond to extreme rays of the corresponding polyhedron.
@@ 4.1 SPO+
Authors in Elmachtoub and Grigas, (2022) define the SPO+ loss function as a convex surrogate loss, which serves as an upper bound on the true regret. The authors derive an informative subgradient for this loss, which can be utilized in any subgradient descent-based approach. Additionally, they present a linear programming formulation for cases where the nominal problem is linear, and the model for estimation is also linear. This surrogate loss is defined as
$\ell_{SPO}(c,m(\omega,x))=$ $\max_{v\in V}\left\{(c-2m(\omega,x))^{\top}v\right\}+2m(\omega,x))^{\top}v^{*}(c)-z^{*}(c).$
@@
$\min_{\omega,\rho}\quad\frac{1}{N}$ $\sum_{i\in[N]}-b^{\top}\rho_{i}+\mathbf{1}^{\top}\theta^{i}+2m(\omega,x^{i})^{\top}v^{*}(c^{i})$ (13a)
s.t. $\,-\rho_{i}^{\top}A+\theta^{i}\geq c^{i}-2m(\omega,x^{i})$ $i\in[N]$ (13b)
$\rho^{i},\theta^{i}\geq 0$ $i\in[N].$ (13c)
Variables $\rho$ and $\theta$ are dual variables associated to constraints (9b) and (9c), respectively. We consider this formulation as a starting point and as a benchmark against which we will compare our methods.
@@ 4.3 Penalization
Based on the ideas of Aboussoror and Mansouri, (2005), we propose the following related formulation: we fix the variables $\gamma^{i}$ in (12) to have all the same fixed value $\kappa$. This parameter $\kappa$ is set before optimization and can be seen as a hyperparameter of the optimization problem. This results in the following model.
$\min_{\omega,\mu,\theta,\delta}\quad$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+\mathbf{1}^{\top}\theta^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)$ (14a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\kappa+\theta^{i}\geq\frac{1}{N}c^{i}$ $\forall i\in[N]$ (14b)
@@
$-\kappa\mathbf{1}+\delta^{i}\leq 0$ $\forall i\in[N]$ (14d)
$\mu^{i}\leq 0$ $\forall i\in[N]$ (14e)
$\delta^{i},\theta^{i}\geq 0$ $\forall i\in[N]$ (14f)
This formulation corresponds to a slice of the formulation (12) which removes all non-linearities of the constraints. Consequently, by adopting this approach, we solve an optimization problem with quadratic non-convex objective and linear constraints.
We remark that this approach results in a problem that is not a traditional penalization (which typically yields relaxations) but rather a restriction of the problem. The name “penalization”, which is used in Aboussoror and Mansouri, (2005), comes from a derivation of (14), which follows a similar approach to the one described in Section 3. The difference lies in penalizing (11h) using $\kappa$ before taking the dual a second time.
@@ 4.4 Alternating direction method
We note that the non-convexities of formulation (12) come from products between the model parameter $\omega$ and dual variables $\gamma$ or $\delta$. Hence, if we fix either the $\omega$ variables, or variables $\gamma$ and $\delta$, we obtain linear programming problems. Specifically, if we fix $\omega$ to $\bar{\omega}$ the resulting LP reads
$\min_{\mu,\theta,\delta,\gamma}\,$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+\mathbf{1}^{\top}\theta^{i}+m(\bar{\omega},x^{i})^{\top}\delta^{i}\right)$ (15a)
s.t. $A^{\top}\mu^{i}+m(\bar{\omega},x^{i})\gamma^{i}+\theta^{i}\geq\frac{1}{N}c^{i}$ $\forall i\in[N]$ (15b)
@@
$-\gamma^{i}\mathbf{1}+\delta^{i}\leq 0$ $\forall i\in[N]$ (15d)
$\mu^{i}\leq 0$ $\forall i\in[N]$ (15e)
$\delta^{i},\gamma^{i},\theta^{i}\geq 0$ $\forall i\in[N].$ (15f)
@@
$\min_{\mu,\theta,\omega}\quad$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+\mathbf{1}^{\top}\theta^{i}+m(\omega,x^{i})^{\top}\bar{\delta}^{i}\right)$ (16a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\bar{\gamma}^{i}+\theta^{i}\geq\frac{1}{N}c^{i}$ $\forall i\in[N]$ (16b)
$\mu^{i}\leq 0,\,\,\theta^{i}\geq 0$ $\forall i\in[N]$ (16c)
@@ 4.5 Regression bounds and valid inequalities
To prevent the solver from generating solutions with large coefficients in the non-convex QCQPs, and since the lower-level optimization problem is invariant to scalings of the objective, we can impose arbitrary bounds to the values of $\omega$. Any bound is valid, but we avoided small numbers to prevent numerical instabilities.
@@
$\sum_{i\in[N]}\left(b^{\top}\mu^{i}+\mathbf{1}^{\top}\theta^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)\geq\frac{1}{N}\sum_{i\in[N]}z^{*}(c^{i})$
This is a dual cut-off constraint. Its left hand side takes the same value as $\frac{1}{N}\sum_{i\in N}(c^{i})^{\top}v^{i}$ in (10), and thus, by optimality of $z^{*}$, the inequality holds. This simple inequality provided considerable improvements in solver’s performance.
@@ 5.1 Computational set-up
We consider an adaptation of the data generation process described in Elmachtoub and Grigas, (2022) and Tang and Khalil, (2024). The training data consists of $\{(x^{i},c^{i})\}_{i=1}^{N}$ generated synthetically the following way.
We consider two families of instances: small instances including values of $N=\{50,100,200\}$ where all methods can be run in moderate running times, and instances with $N=1000$ to test the scalability of our approach. Each dataset is separated in 70% for training and 30% for testing. Feature vectors are generated by sampling them from a standard normal distribution (mean zero and standard deviation equals to 1). We generate cost vectors by first generating the parameters $\omega$ of the model –representing the true underlying model– and then using the following formula (see Elmachtoub and Grigas, (2022) and Tang and Khalil, (2024)):
@@
We remark that the exact reformulation requires solving a challenging non-convex problem, and that the underlying problem is NP-hard even for one observation (as established in Section 2); these are the main reasons why we focus on moderate instance sizes in this work. To provide some perspective, in Table 1 we provide the number of variables and constraints (linear and non-convex) that the non-convex quadratic problem has.
Variables $N\left(\#\mbox{Nodes}+2\cdot\#\mbox{Edges}+1\right)+K\#\mbox{Edges}$
Constraints Linear (excluding var. bounds) $N(\#\mbox{Nodes}+\#\mbox{Edges})$
Non-convex $N\#\mbox{Edges}$
Table 1: Description of the size of the exact reformulation.
@@
      Use the Alternating Method (Algorithm 2) with the solution produced by LS as starting point
@@
Figure 2 shows that $\epsilon=1$ (size of the neighborhood) and $T=20$ (samples per iteration) provide the best trade-off between test and training performance; we thus set these values for experiments involving bipartite matching. In the case of shortest path, we determined that $\epsilon=0.1$ and $T=20$ yields the best results.
@@
We implemented all the aforementioned routines using Python 3.10. All non-convex quadratic models were solved using Gurobi 10.0.3 (Gurobi Optimization, LLC,, 2023). All experiments were run single-threaded on a Linux machine with an Intel Xeon Silver 4210 2.2G CPU and 128 GB RAM.

### v5

@@ 2.1 Membership in NP
Corollary 3.
SIMPLE-REGRET is NP-complete.
Corollary 4.
SIMPLE-REGRET is not $\Sigma_{2}^{P}$-hard, unless the polynomial hierarchy collapses at the second level.
As a final remark on this subsection, we note that the previous proof strategy can be used to show two facts in the optimization context of (4) when $m(\omega,x)=\omega x$.
Corollary 5.
Consider (4) when $m(\omega,x)=\omega x$. The regret function (2) only has a finite number of values, and, furthermore, the minimum regret (4) is always attained.
The piecewise constant nature of the regret (2) is known in the literature (e.g., poganvcic2019differentiation ; demirovic2020dynamic ). The fact that it only attains a finite number of values in our setting, although perhaps expected, is not entirely direct. Also, recall that some bilevel optimization problems do not attain their optimal values.
The proof of Corollary 5 follows from the fact that, once we fix the optimal faces of each lower-level problem, the regret is fixed. For completeness, we provide the proof of Corollary 5 in Appendix A.
2.2 Polynomial-time solvable cases
In this subsection, we show that in many cases, determining if the regret is 0 can be done in polynomial time. This is in line with the polynomial solvability of checking if zero loss can be achievable in empirical risk minimization in multiple cases. For example, checking if zero loss can be achieved in 0-1 loss minimization (i.e., if the data is separable) can be solved in polynomial time: it amounts to finding a separating hyperplane among the two classes. However, we note that, perhaps counterintuitively, having zero regret is not equivalent to determining if there exists $\omega$ such that
$\omega x^{i}=c^{i}\qquad\forall i\in[N].$
We illustrate this in the following example.
Example 6.
Consider the polytope $V=[0,1]^{2}$ and the following two observations
$(x^{1},c^{1})=\left(1,\begin{pmatrix}-1,-2\end{pmatrix}^{\top},\right)\,(x^{2},c^{2})=\left(-1,\begin{pmatrix}1,1\end{pmatrix}^{\top}\right)$.
It is not hard to see that
$\arg\min\{c^{1}{}^{\top}v\,:\,v\in[0,1]^{2}\}=\{(1,1)^{\top}\}\quad\text{and}\quad\arg\min\{c^{2}{}^{\top}v\,:\,v\in[0,1]^{2}\}=\{(0,0)^{\top}\}$
If we take, for example, $\omega=(-1,-1)^{\top}$, we have that
$\arg\min\{(\omega x^{1})^{\top}v\,:\,v\in[0,1]^{2}\}=\{(1,1)^{\top}\}\text{ and }\arg\min\{(\omega x^{2})^{\top}v\,:\,v\in[0,1]^{2}\}=\{(0,0)^{\top}\}$
And from this, we can deduce that $\omega$ yields zero regret. However, it is not hard to see that there is no $\omega$ such that $\omega x^{i}=c^{i},\,i=1,2,$ as $(\omega x^{1},\omega x^{2})$ are always collinear in this example, but $(c^{1},c^{2})$ are linearly independent.
Our polynomial-time solvability result uses the following assumption.
Assumption 7.
The input data $(c^{i},x^{i})_{i=1}^{N}$ and the polytope $V$ are such that $\arg\min\{c^{i}{}^{\top}v\,:\,v\in V\}$ is a singleton, for all $i\in[N]$.
Assumption 7 may seem restrictive, but it can be expected for real data to satisfy it; if the $c^{i}$ are drawn from a non-atomic distribution, for example, Assumption 7 is satisfied with probability 1. Additionally, note that this does not imply that the follower will have a unique solution in general: it can still be that the optimal face for $\omega x^{i}$ over $V$ is not a singleton. Finally, note that Assumption 7 can be checked in polynomial time.
We are now ready to describe a polynomial time solvable case for (4).
Theorem 8.
If the input for SIMPLE-REGRET is restricted to $M=0$ (i.e., determining if there is a solution with zero regret) and the data $(c^{i},x^{i})_{i=1}^{N}$ and the polytope $V$ satisfy Assumption 7, then the problem can be solved in polynomial time.
We provide the proof of Theorem 8 in Appendix B. As a last comment in this subsection, we conjecture that, without Assumption 7, the problem of determining if there is a solution with zero regret becomes NP-hard. This is somewhat counterintuitive since some NP-hard loss minimization problems (as 0-1 loss minimization) become easy when the question is whether zero loss is achievable or not. However, we conjecture this based on the discussion on Example 6: one can have zero regret with data that cannot be fit perfectly in the traditional sense, and this can be obtained with a non-trivial “alignment" of the optimal faces $V^{*}(\omega x^{i})$.
@@ 3 A non-convex quadratic reformulation
While NP-hardness of (4) makes it unlikely for us to find a worst-case efficient method for solving it, we can still hope to find methods with good practical performance. In this section, we will apply duality arguments, supported by the assumption that $V$ in problem (1) is non-empty and bounded, in order to obtain a more manageable formulation of the problem.
As before, we assume (1) has the following form:
$\min_{v}\quad$ $c^{\top}v\quad\text{s.t.}\quad Av\geq b$ (7a)
@@
$\quad\quad\quad\quad\text{s.t.}\quad\quad A\tilde{v}^{i}\geq b$
@@ 3.1 Duality arguments
A common approach to solving optimistic bilevel problems involving convex lower-level problems is to reformulate them by replacing the lower-level problems with their optimality (Karush-Kuhn-Tucker or KKT) conditions (see kleinert2021survey ). In this subsection, we follow the same type of argument twice to achieve a single-level reformulation of the pessimistic bilevel problem (8).
@@
$\quad A^{\top}\rho^{i}=m(\omega,x^{i})$ $\forall i\in[N]$ (9c)
$\quad\rho^{i}\geq 0$ $\forall i\in[N]$ (9d)
$\quad m(\omega,x^{i})^{\top}v^{i}\leq b^{\top}\rho^{i}$ $\forall i\in[N]$ (9e)
In this formulation (9b) imposes primal feasibility, (9c)-(9d) impose dual feasibility, and (9e) imposes strong duality. We note that strong duality is typically written as an equality constraint; however, the $\geq$ inequality always holds due to weak duality. We purposely opted for imposing strong duality directly instead of complementary slackness, as the latter yields non-linear inequalities on $(v,\rho)$.
@@
$\min_{\omega,\mu,\delta,\gamma}\,$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)$ (10a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\gamma^{i}=\frac{1}{N}c^{i}$ $\forall i\in[N]$ (10b)
@@
$\mu^{i}\leq 0,\,\gamma^{i}\geq 0$ $\forall i\in[N]$ (10d)
@@ 3.2 Shortest path as a bounded linear program
In this work, and motivated by elmachtoub2022smart , we consider the shortest path problem with unknown cost vectors. It is well known that this problem can be formulated as a linear program using a totally unimodular constraint matrix. However, the feasible region may not be bounded, as the underlying graph may have negative cycles, which correspond to extreme rays of the corresponding polyhedron.
@@ 4.1 SPO+
Authors in elmachtoub2022smart define the SPO+ loss function as a convex surrogate loss, which serves as an upper bound on the true regret. The authors derive an informative subgradient for this loss, which can be utilized in any subgradient descent-based approach. Additionally, they present a linear programming formulation for cases where the nominal problem is linear, and the model for estimation is also linear. This surrogate loss is defined as
$\ell_{SPO}(c,m(\omega,x))=$ $\max_{v\in V}\left\{(c-2m(\omega,x))^{\top}v\right\}+2m(\omega,x)^{\top}v^{*}(c)-z^{*}(c).$
@@
$\min_{\omega,\rho}\quad\frac{1}{N}$ $\sum_{i\in[N]}\left(-b^{\top}\rho_{i}+2m(\omega,x^{i})^{\top}v^{*}(c^{i})\right)$ (11a)
s.t. $\,-\rho_{i}^{\top}A=c^{i}-2m(\omega,x^{i})$ $i\in[N]$ (11b)
$\rho^{i}\geq 0$ $i\in[N].$ (11c)
Variables $\rho$ are dual variables associated to constraints $Av\geq b$ of the nominal problem (7). We consider this formulation as a starting point and as a benchmark against which we will compare our methods.
@@ 4.3 Penalization
Based on the ideas of aboussoror2005weak , we propose the following related formulation: we fix the variables $\gamma^{i}$ in (10) to have all the same fixed value $\kappa$. This parameter $\kappa$ is set before optimization and can be seen as a hyperparameter of the optimization problem. This results in the following model.
$\min_{\omega,\mu,\delta}\quad$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)$ (12a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\kappa=\frac{1}{N}c^{i}$ $\forall i\in[N]$ (12b)
@@
$\mu^{i}\leq 0$ $\forall i\in[N]$ (12d)
This formulation corresponds to a slice of the formulation (10) which removes all non-linearities of the constraints. Consequently, by adopting this approach, we solve an optimization problem with a quadratic non-convex objective and linear constraints.
We remark that this approach results in a problem that is not a traditional penalization (which typically yields relaxations) but rather a restriction of the problem. The name “penalization”, which is used in aboussoror2005weak , comes from a derivation of (12), which follows a similar approach to the one described in Section 3. The difference lies in penalizing (9e) using $\kappa$ before taking the dual a second time.
@@ 4.4 Alternating direction method
We note that the non-convexities of formulation (10) come from products between the model parameter $\omega$ and dual variables $\gamma$ or $\delta$. Hence, if we fix either the $\omega$ variables or the variables $\gamma$ and $\delta$, we obtain linear programming problems. Specifically, if we fix $\omega$ to $\bar{\omega}$ the resulting LP reads
$\min_{\mu,\delta,\gamma}\,$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+m(\bar{\omega},x^{i})^{\top}\delta^{i}\right)$ (13a)
s.t. $A^{\top}\mu^{i}+m(\bar{\omega},x^{i})\gamma^{i}=\frac{1}{N}c^{i}$ $\forall i\in[N]$ (13b)
@@
$\mu^{i}\leq 0$ $\forall i\in[N]$ (13d)
$\gamma^{i}\geq 0$ $\forall i\in[N].$ (13e)
@@
$\min_{\mu,\omega}\quad$ $\sum_{i\in[N]}\left(b^{\top}\mu^{i}+m(\omega,x^{i})^{\top}\bar{\delta}^{i}\right)$ (14a)
s.t. $A^{\top}\mu^{i}+m(\omega,x^{i})\bar{\gamma}^{i}=\frac{1}{N}c^{i}$ $\forall i\in[N]$ (14b)
$\mu^{i}\leq 0$ $\forall i\in[N]$ (14c)
@@ 4.5 Regression bounds and valid inequalities
To prevent the solver from generating solutions with large coefficients in the non-convex QCQPs, and since the lower-level optimization problem is invariant to scalings of the objective, we can impose arbitrary bounds on the values of $\omega$. Any bound is valid, but we avoided small numbers to prevent numerical instabilities.
@@
$\sum_{i\in[N]}\left(b^{\top}\mu^{i}+m(\omega,x^{i})^{\top}\delta^{i}\right)\geq\frac{1}{N}\sum_{i\in[N]}z^{*}(c^{i})$
This is a dual cut-off constraint. Its left hand side takes the same value as $\frac{1}{N}\sum_{i\in N}(c^{i})^{\top}v^{i}$ in (8), and thus, by optimality of $z^{*}$, the inequality holds. This simple inequality provided considerable improvements in the solver’s performance.
@@ 5.1 Computational set-up
We consider an adaptation of the data generation process described in elmachtoub2022smart and tang2022pyepo . The training data consists of $\{(x^{i},c^{i})\}_{i=1}^{N}$ generated synthetically the following way.
We consider two families of instances: small instances including values of $N=\{50,100,200\}$ where all methods can be run in moderate running times, and instances with $N=1000$ to test the scalability of our approach. Each dataset is separated into 70% for training and 30% for testing. Feature vectors are generated by sampling them from a standard normal distribution (mean zero and standard deviation equal to 1). We generate cost vectors by first generating the parameters $\omega$ of the model –representing the true underlying model– and then using the following formula (see elmachtoub2022smart and tang2022pyepo ):
@@
We remark that the exact reformulation requires solving a challenging non-convex problem, and that the underlying problem is NP-complete; these are the main reasons why we focus on moderate instance sizes in this work. To provide some perspective, the reformulation involves $N\cdot\left(\#\mbox{Nodes}+2\cdot\#\mbox{Edges}+1\right)+K\cdot\#\mbox{Edges}$ variables and $N\cdot(\#\mbox{Nodes}+\#\mbox{Edges})$ constraints, from which $N\cdot\#\mbox{Edges}$ are quadratic non-convex constraints.
@@
      Use the Alternating Method (Algorithm 2) with the solution produced by LS as the starting point
@@
Figure 2 shows that $\epsilon=1$ (size of the neighborhood) and $T=20$ (samples per iteration) provide the best trade-off between test and training performance; we thus set these values for experiments involving bipartite matching. In the case of the shortest path, we determined that $\epsilon=0.1$ and $T=20$ yields the best results.
@@
We implemented all the aforementioned routines using Python 3.10. All non-convex quadratic models were solved using Gurobi 10.0.3 (gurobi, ). All experiments were run single-threaded on a Linux machine with an Intel Xeon Silver 4210 2.2G CPU and 128 GB RAM.
Repository
The codes and instances considered in this article can be found in the repository in the following link.